### v2

@@ Appendix B Significance analysis
Figure 4. Visualization of significance analysis across models. (a) shows the pairwise-bootstrap significance matrix, while (b) summarizes how many significant differences each model has relative to others.
@@
Appendix C Literature Review
@@
C.1. XBRL Tagging benchmark
@@
FiNER systematically benchmarked several neural architectures on the finer-139 dataset to address numeric-heavy XBRL tagging (Loukas et al., 2022). The initial experiments showed that standard BERT underperforms due to subword fragmentation, then the authors introduced pseudo-token strategies replacing numerals with [NUM] or [SHAPE] tokens to stabilize label assignment across fragmented numeric spans. These strategies, combined with domain-specific pretraining on SEC-BERT, significantly improved tagging performance, reaching 82.1 micro-F1 without the need for computationally expensive CRF layers. Their experiments demonstrated that subword-aware models with numeric-aware pseudo-tokens outperform both word-level BiLSTMs and vanilla BERT, particularly in numeric-heavy contexts, and avoid nonsensical label sequences. FNXL extended this benchmarking paradigm to a much larger label space of 2,794 US-GAAP tags, reframing the task as an extreme classification problem (Sharma et al., 2023). They compared the FiNER sequence labeling approach with a two-step pipeline that first identifies numeric spans and then assigns labels using AttentionXML. While FiNER achieved stronger micro-F1 (75.84), reflecting better performance on frequent tags, AttentionXML outperformed FiNER in macro-F1 (47.54), highlighting its strength in predicting infrequent, tail-end labels. FNXL further evaluated both models under a Hits@k setting, confirming that label recommendations from the AttentionXML pipeline could substantially reduce manual effort and maintain high inter-annotator agreement. Together, these benchmarks reveal the need for context-aware reasoning and label-ranking mechanisms in realistic XBRL tagging scenarios.
C.2. XBRL Tagging Methods
@@
The previous studies also explored the approaches address scalability, semantic ambiguity, and reasoning gaps in XBRL tagging to improve the performance. Saini et al. (Saini et al., 2021) proposed GalaXC is a graph-based extreme classification framework that jointly learns over document-label graphs with per-label attention across multi-hop neighborhoods. By integrating label metadata and transitive label correlations, GalaXC outperformed leading deep classifiers by up to 18% in micro-F1 on standard benchmarks and achieved 25% gains in warm-start scenarios where partial labels are available. Moreover, Wang et al. (Wang, 2023) addressed the practical challenge of custom tag standardization through a semantic similarity pipeline that leverages TF-IDF, Word2Vec, and FinBERT embeddings. Although unsupervised, the method was tested across nearly 200,000 custom tags from SEC filings between 2009 and 2022, and showed strong alignment performance, with vector-based mappings identifying viable standard tag candidates for a substantial proportion of non-compliant elements—offering a low-cost, interpretable solution for downstream financial analysis. Shifting focus from classification to comprehension, XBRL-Agent evaluated the capabilities of large language models to reason over full XBRL reports (Han et al., 2024). The authors introduced two task types—domain taxonomy understanding and numeric reasoning and found that base LLMs often hallucinated or misinterpreted financial content. To overcome these issues, XBRL-Agent incorporated retrieval-augmented generation (RAG) and symbolic calculators within an LLM-agent framework. The enhanced system achieved a 17% accuracy gain on domain query tasks and a 42% boost on numeric reasoning queries compared to base LLMs, validating the utility of modular tool augmentation. These improvements enabled reliable multi-step reasoning over complex disclosures such as debt instruments and derivative gains, which are difficult to capture using span-level classifiers. Collectively, these works broaden the methodological landscape of XBRL tagging from graph-based label propagation and embedding-based normalization to LLM-driven report analysis and point to a hybrid future where structured priors and reasoning tools jointly support accurate, scalable financial information extraction.
C.3. Financial Evaluation Benchmarks
@@
In parallel to XBRL-specific advances, the financial NLP community has developed comprehensive benchmarks to assess broader capabilities in information extraction, numerical reasoning, and document understanding. FiNER-ORD (Shah et al., 2023) introduced a high-quality, domain-specific NER dataset annotated over financial news, emphasizing general entity types like persons, organizations, and locations. While not numerically focused like FiNER-139, it highlights the lexical diversity of financial discourse and establishes a strong baseline for testing pretrained and zero-shot LLMs in real-world financial NER scenarios. FinQA (Chen et al., 2021) pushed toward explainable QA by pairing expert-written questions with annotated multi-step reasoning programs derived from earnings reports. ConvFinQA (Chen et al., 2022) extended this challenge to conversational contexts, simulating real-world question flows over sequential financial queries. TAT-QA (Zhu et al., 2021) focused on hybrid tabular-text reasoning and required models to align cell values and document narratives, often involving aggregation, comparison, and unit-scale interpretation. Pixiu (Xie et al., 2023) introduced a broader evaluation framework by releasing FinMA, a financial LLM instruction-tuned across five tasks, and assessing it on a new benchmark covering sentiment classification, QA, summarization, NER, and stock prediction. BizBench (Koncel-Kedziorski et al., 2023) framed financial QA as program synthesis over realistic, multi-modal contexts, integrating reasoning, code generation, and domain knowledge into a single evaluation pyramid. While these benchmarks highlight the growing ability of models to integrate structured and unstructured financial data, they overlook taxonomy-driven fact alignment and do not support the structured output formats required for XBRL tagging.
Appendix D Data and Ticker Information
D.1. Overall Data Annotation
Different from the filtered dataset described in Section 3.3, we construct an overall annotation dataset based directly on the unfiltered raw data. As summarized in Table 11, the dataset consists of both sentence-level and table-level instances drawn from the narrative and tabular content of financial reports. In total, we annotated 15,986 sentence instances and 12,801 table instances, including 7,768 positive and 8,218 negative sentences, as well as 8,709 positive and 4,092 negative tables. On average, textual sentences contain 80.91 tokens, while table sequences are considerably longer, averaging 1,212.42 tokens. Across all data, we identified 261,457 numerical entities linked to 3,953 unique US-GAAP concepts.
Table 11. Statistical information for the original dataset in our benchmark. Tokens are calculated using the “cl100k_base” tokenizer ($\pm$ standard deviation).
Item Sentence Table
Positive instances 7,768 8,709
Negative instances 8,218 4,092
Avg. Tokens/S $80.91\pm 63.62$ $1212.42\pm 1421.76$
Avg. Entities/S $1.24\pm 1.82$ $18.87\pm 37.16$
Avg. Concepts/S $1.24\pm 1.82$ $18.87\pm 37.16$
Total Entities 261,457
Unique Concepts 3,953
D.2. The Statistics of the Tickers
@@
Figure 5, Figure 6, and Figure 7 illustrate the distribution of 142 tickers across industry sectors, market capitalization categories, and geographic regions, highlighting the diversity of our collected financial reports. Considering that practical XBRL tagging practices often vary by industry, company size, and legal jurisdiction, we curated a diverse set of companies covering all 11 major industry sectors. The sample maintains a balanced distribution across firm sizes, following market capitalization categories: micro-cap (¡\$300M), small-cap (\$300M–\$2B), mid-cap (\$2B–\$10B), large-cap (\$10B–\$200B), and mega-cap (¿\$200B). In addition, we incorporated firms from over 30 states and international jurisdictions to capture regional differences in reporting and tagging conventions. This diversity ensures that the benchmark reflects realistic heterogeneity observed in financial disclosures across sectors, scales, and regulatory environments.
Figure 5. Distribution of Tickers by Industry Sector.
Figure 6. Distribution of Tickers by Market Cap Size.
Figure 7. Distribution of Tickers by Geography.
Appendix E Benchmark Granularity Justification
@@
Appendix F Validation Guideline
F.1. Task Definition
@@ F.2. Validation Procedure
Table 12. Entity type definitions used for validation.
@@ F.3. Important Instructions
Appendix G Evaluation Metrics
To provide a fair evaluation of overall benchmark performance, we adopt a set of metrics, focusing primarily on macro-level and micro-level evaluation strategies inspired by the previous work (Sharma et al., 2023). Macro-level evaluation computes precision, recall, and F1 scores independently for each BIO-concept label derived from the US-GAAP taxonomy, and then averages them without weighting. This ensures that each concept, including rare or infrequent ones, contributes equally to the final score, making it especially suitable for domains with skewed label distributions. In contrast, micro-level evaluation aggregates token-level true positives, false positives, and false negatives across all labels before computing precision, recall, and F1. This approach emphasizes the model’s overall tagging accuracy by treating every token equally and thus better reflects performance on frequent concepts. Together, these two metrics provide a balanced view of both per-concept performance and overall tagging quality.
@@ Appendix G Evaluation Metrics
(5) $\text{Precision}=\frac{|\mathcal{G}\cap\mathcal{P}|}{|\mathcal{P}|}$
(6) $\text{Recall}=\frac{|\mathcal{G}\cap\mathcal{P}|}{|\mathcal{G}|}$
(7) $\text{F1}=\frac{2\times\text{Precision}\times\text{Recall}}{\text{Precision}+\text{Recall}}$
@@
(8) $\text{Accuracy}=\frac{1}{N}\sum_{i=1}^{N}\delta(\hat{c}_{i},c_{i}^{*})$
@@
Appendix H Evaluation Models Details

### v3

@@ Appendix A Significance analysis
Figure 4: Visualization of significance analysis across models. (a) shows the pairwise-bootstrap significance matrix, while (b) summarizes how many significant differences each model has relative to others.
@@
Appendix B Literature Review
@@
B.1 XBRL Tagging benchmark
@@
FiNER systematically benchmarked several neural architectures on the finer-139 dataset to address numeric-heavy XBRL tagging loukas2022finer. The initial experiments showed that standard BERT underperforms due to subword fragmentation, then the authors introduced pseudo-token strategies replacing numerals with [NUM] or [SHAPE] tokens to stabilize label assignment across fragmented numeric spans. These strategies, combined with domain-specific pretraining on SEC-BERT, significantly improved tagging performance, reaching 82.1 micro-F1 without the need for computationally expensive CRF layers. Their experiments demonstrated that subword-aware models with numeric-aware pseudo-tokens outperform both word-level BiLSTMs and vanilla BERT, particularly in numeric-heavy contexts, and avoid nonsensical label sequences. FNXL extended this benchmarking paradigm to a much larger label space of 2,794 US-GAAP tags, reframing the task as an extreme classification problem sharma2023financial. They compared the FiNER sequence labeling approach with a two-step pipeline that first identifies numeric spans and then assigns labels using AttentionXML. While FiNER achieved stronger micro-F1 (75.84), reflecting better performance on frequent tags, AttentionXML outperformed FiNER in macro-F1 (47.54), highlighting its strength in predicting infrequent, tail-end labels. FNXL further evaluated both models under a Hits@k setting, confirming that label recommendations from the AttentionXML pipeline could substantially reduce manual effort and maintain high inter-annotator agreement. Together, these benchmarks reveal the need for context-aware reasoning and label-ranking mechanisms in realistic XBRL tagging scenarios.
B.2 XBRL Tagging Methods
@@
The previous studies also explored approaches to address scalability, semantic ambiguity, and reasoning gaps in XBRL tagging to improve performance. Saini et al. saini2021galaxc proposed GalaXC is a graph-based extreme classification framework that jointly learns over document-label graphs with per-label attention across multi-hop neighborhoods. By integrating label metadata and transitive label correlations, GalaXC outperformed leading deep classifiers by up to 18% in micro-F1 on standard benchmarks and achieved 25% gains in warm-start scenarios where partial labels are available. Moreover, Wang et al. wang2023standardizing addressed the practical challenge of custom tag standardization through a semantic similarity pipeline that leverages TF-IDF, Word2Vec, and FinBERT embeddings. Although unsupervised, the method was tested across nearly 200,000 custom tags from SEC filings between 2009 and 2022, and showed strong alignment performance, with vector-based mappings identifying viable standard tag candidates for a substantial proportion of non-compliant elements, offering a low-cost and interpretable solution for downstream financial analysis. Shifting focus from classification to comprehension, XBRL-Agent evaluated the capabilities of large language models to reason over full XBRL reports han2024xbrlagent. The authors introduced two task types, domain taxonomy understanding and numeric reasoning, and found that base LLMs often hallucinated or misinterpreted financial content. To overcome these issues, XBRL-Agent incorporated retrieval-augmented generation (RAG) and symbolic calculators within an LLM-agent framework. The enhanced system achieved a 17% accuracy gain on domain query tasks and a 42% boost on numeric reasoning queries compared to base LLMs, validating the utility of modular tool augmentation. These improvements enabled reliable multi-step reasoning over complex disclosures such as debt instruments and derivative gains, which are difficult to capture using span-level classifiers. Collectively, these works broaden the methodological landscape of XBRL tagging from graph-based label propagation and embedding-based normalization to LLM-driven report analysis and point to a hybrid future where structured priors and reasoning tools jointly support accurate, scalable financial information extraction.
B.3 Financial Evaluation Benchmarks
@@
In parallel to XBRL-specific advances, the financial NLP community has developed comprehensive benchmarks to assess broader capabilities in information extraction, numerical reasoning, and document understanding. FiNER-ORD shah2023finer introduced a high-quality, domain-specific NER dataset annotated over financial news, emphasizing general entity types like persons, organizations, and locations. While not numerically focused like FiNER-139, it highlights the lexical diversity of financial discourse and establishes a strong baseline for testing pretrained and zero-shot LLMs in real-world financial NER scenarios. FinQA chen2021finqa pushed toward explainable QA by pairing expert-written questions with annotated multi-step reasoning programs derived from earnings reports. ConvFinQA chen2022convfinqa extended this challenge to conversational contexts, simulating real-world question flows over sequential financial queries. TAT-QA zhu2021tat focused on hybrid tabular-text reasoning and required models to align cell values and document narratives, often involving aggregation, comparison, and unit-scale interpretation. Pixiu xie2023pixiu introduced a broader evaluation framework by releasing FinMA, a financial LLM instruction-tuned across five tasks, and assessing it on a new benchmark covering sentiment classification, QA, summarization, NER, and stock prediction. BizBench koncel2023bizbench framed financial QA as program synthesis over realistic, multi-modal contexts, integrating reasoning, code generation, and domain knowledge into a single evaluation pyramid. While these benchmarks highlight the growing ability of models to integrate structured and unstructured financial data, they overlook taxonomy-driven fact alignment and do not support the structured output formats required for XBRL tagging.
Appendix C Data and Ticker Information
C.1 The Statistics of the Tickers
@@
Figure 5, Figure 6, and Figure 7 illustrate the distribution of 142 tickers across industry sectors, market capitalization categories, and geographic regions, highlighting the diversity of our collected financial reports. Considering that practical XBRL tagging practices often vary by industry, company size, and legal jurisdiction, we curated a diverse set of companies covering all 11 major industry sectors. The sample maintains a balanced distribution across firm sizes, following market capitalization categories: micro-cap (<\$300M), small-cap (\$300M–\$2B), mid-cap (\$2B–\$10B), large-cap (\$10B–\$200B), and mega-cap (>\$200B). In addition, we incorporated firms from over 30 states and international jurisdictions to capture regional differences in reporting and tagging conventions. This diversity ensures that the benchmark reflects realistic heterogeneity observed in financial disclosures across sectors, scales, and regulatory environments.
Figure 5: Distribution of Tickers by Industry Sector.
Figure 6: Distribution of Tickers by Market Cap Size.
Figure 7: Distribution of Tickers by Geography.
C.2 Data Granularity Justification
@@
C.3 Automated Data Annotation
Algorithm 1 converts Inline XBRL (iXBRL) filings into structured instances for our benchmark. The procedure operates over both narrative sentences and tables, and assigns each instance a positive or negative label based on whether it contains at least one validated financial entity.
We begin by iterating through all disclosure instances in the corpus and extracting their plain-text content. Very short fragments (length $\leq 20$) and duplicated passages are discarded using a global text set, ensuring that the benchmark contains unique and meaningful contexts.
For each remaining instance, the algorithm traverses all atomic iXBRL elements. From each element, we retrieve the numerical value $v$, its associated concept $c$ from the US-GAAP taxonomy, and the declared data type $\tau$. We retain only elements whose concept is in $\mathcal{T}$ and whose type belongs to the predefined valid numeric set $\mathcal{L}$. Restricting extraction to leaf-level elements avoids double counting values that appear in nested XML structures. A local deduplication set further prevents repeated annotation of identical $(v,c)$ pairs within the same instance.
If at least one valid entity is identified, the instance is recorded as a positive example together with its entity list; otherwise, it is stored as a negative example. Negative instances are intentionally preserved, as they reflect realistic disclosure text that models must learn to ignore. Overall, this process yields a corpus that is faithful to regulatory metadata while remaining structurally clean and suitable for downstream evaluation.
Input: Raw 10-K reports (iXBRL) $\mathcal{D}$; US-GAAP taxonomy $\mathcal{T}$; Set of valid data types $\mathcal{L}$
Output: Positive instances $\mathcal{P}$; Negative instances $\mathcal{N}$
Initialize: $\mathcal{P}\leftarrow\emptyset$, $\mathcal{N}\leftarrow\emptyset$, Global unique text set $\mathcal{S}_{text}\leftarrow\emptyset$;
foreach instance (sentence or table) $i\in\mathcal{D}$ do
 $text\leftarrow$ Extract plain text content from instance $i$;
 if length of $text\leq 20$ or $text\in\mathcal{S}_{text}$ then
    skip to the next instance;
 $E\leftarrow\emptyset$ // List of validated entities in $i$;
 $S_{fact}\leftarrow\emptyset$ // Local fact deduplication set;
 foreach atomic iXBRL element $\epsilon$ within instance $i$ do
      Retrieve value $v$, financial concept $c$, and data type $\tau$ from $\epsilon$;
    if $c\in\mathcal{T}$ and $\tau\in\mathcal{L}$ and $v$ is not empty then
       if $(v,c)\notin S_{fact}$ then
            Add $(v,\tau,c)$ to the entity list $E$;
            Add $(v,c)$ to the local deduplication set $S_{fact}$;
 if $E$ is not empty then
      Add $(text,E)$ to the positive set $\mathcal{P}$;
 else
      Add $(text,\emptyset)$ to the negative set $\mathcal{N}$;
  Add texttext to the global set $\mathcal{S}_{text}$;
return $\mathcal{P},\mathcal{N}$;
Algorithm 1 Automated Instance Annotation
Appendix D Annotation Data Quality Control
D.1 Validation Data Filtering
To establish a reliable upper bound on annotation quality, we construct a strategic audit subset rather than validating the full corpus. Algorithm 2 formalizes the sampling process as a greedy minimal-cover problem over observed US-GAAP concepts.
The algorithm takes as input the set of positive instances $\mathcal{P}$, together with the set of all taxonomy concepts observed in the corpus, $\mathcal{C}_{all}$. The goal is to produce a compact subset $\mathcal{P}^{*}$ that is both diverse and concept-complete.
Seeding for company diversity. We first iterate over companies and select up to $K$ instances per company. This prevents the audit set from being dominated by firms with longer filings or more frequent disclosures. After each selection, we update the set of uncovered concepts $\mathcal{C}_{rem}$ by removing concepts already represented among the seeds.
Greedy minimal cover. We then repeatedly add instances that contribute maximal marginal concept coverage. For each candidate instance, we compute the number of uncovered concepts it contains and select the instance with the largest score. When ties occur, we prefer shorter instances to reduce annotation effort while retaining conceptual value. After adding each instance, we update $\mathcal{C}_{rem}$ accordingly and continue until either all concepts are covered or no additional candidates remain.
Resulting audit subset. This procedure yields a compact subset that balances two competing objectives: coverage of the observed taxonomy and representation across companies. The resulting $\mathcal{P}^{*}$ therefore serves as a stress-test sample that is both manageable for human review and representative of the broader benchmark space.
Input: Positive instances $\mathcal{P}$ from Algorithm 1; Observed taxonomy concepts $\mathcal{C}_{all}$; Company cap $K=10$
Output: Strategic audit subset $\mathcal{P}^{*}$
Initialize: $\mathcal{P}^{*}\leftarrow\emptyset$;
Initialize: $\mathcal{C}_{rem}\leftarrow\mathcal{C}_{all}$ // Set of concepts yet to be covered;
Step 1: Diversity-Preserving Seeding
foreach unique company in $\mathcal{P}$ do
   Select up to $K$ instances belonging to this company and add to $\mathcal{P}^{*}$;
   Update $\mathcal{C}_{rem}$ by removing concepts covered by these seed instances;
Step 2: Greedy Minimal Cover
while $\mathcal{C}_{rem}$ is not empty and candidates remain in $\mathcal{P}$ do
 foreach candidate instance $i\in\mathcal{P}\setminus\mathcal{P}^{*}$ do
      Calculate $Score(i)=|\text{Concepts in }i\cap\mathcal{C}_{rem}|$;
  Select instance $i^{*}$ with the highest $Score(i)$;
 if multiple instances have the same $Score(i^{*})$ then
      Break ties by selecting the instance with the shortest text length;
  Add $i^{*}$ to $\mathcal{P}^{*}$ and update $\mathcal{C}_{rem}$ by removing newly covered concepts;
return $\mathcal{P}^{*}$;
Algorithm 2 Strategic Audit Set Sampling
D.2 Annotator Demography
To ensure reliable verification of financial entities and taxonomy mappings, all audits were conducted by annotators with prior exposure to XBRL-based reporting and financial statement analysis. The review team consisted of two junior analysts and one senior adjudicator.
The two junior analysts are graduate-level researchers with training in accounting and financial data analytics. Both have received formal instruction on US-GAAP concepts, XBRL tagging conventions, and common disclosure practices in SEC filings. They have prior experience working with financial datasets in academic or industry projects, including tasks involving concept normalization, document review, and regulatory text interpretation. Before participating in the audit, each analyst completed a guided calibration session using our Verification Guideline to ensure consistent treatment of numerical entities, metadata attributes, and taxonomy semantics.
The senior adjudicator has over a decade of experience in financial reporting analysis and structured disclosure systems. Their professional background includes work with XBRL instances, taxonomy structure, and validation procedures used in regulatory environments. In addition to overseeing the audit workflow, the senior adjudicator resolved disagreements, provided clarification on ambiguous tagging scenarios, and ensured alignment with US-GAAP definitions and XBRL best practices.
Together, this composition provides both annotation robustness and domain oversight: junior analysts contribute detailed review capacity, while the senior adjudicator ensures conceptual correctness and methodological consistency across the audit.
All annotators were members of the research team and were invited to participate as part of their normal research training and collaborative project responsibilities. Participation was voluntary, and no separate monetary compensation was provided. Because the work involved publicly available financial disclosures and did not require annotators to process personal or sensitive data, the review posed minimal risk to participants.
D.3 Validation Guideline
D.3.1 Task Definition
@@ D.3.2 Validation Procedure
Table 10: Entity type definitions used for validation.
@@ D.3.3 Important Instructions
D.4 Agreement Computation
We report both raw agreement and Cohen’s $\kappa$ to quantify inter-annotator reliability.
Let $N$ denote the total number of instances, and let $n_{yy}$ be the number of cases on which both annotators assign the same label $y$. The raw agreement is
$P_{o}=\frac{1}{N}\sum_{y}n_{yy}.$ (5)
To adjust for agreement occurring by chance, we compute Cohen’s $\kappa$. Let $p_{1y}$ and $p_{2y}$ denote the marginal probabilities that annotator 1 and annotator 2 assign label $y$, respectively. The expected chance agreement is
$P_{e}=\sum_{y}p_{1y}\,p_{2y},$ (6)
and Cohen’s $\kappa$ is
$\kappa=\frac{P_{o}-P_{e}}{1-P_{e}}.$ (7)
Following landis1977measurement, our $\kappa=0.81$ indicates substantial agreement.
Appendix E Evaluation Metrics
To provide a fair evaluation of overall benchmark performance, we adopt a set of metrics, focusing primarily on macro-level and micro-level evaluation strategies inspired by the previous work sharma2023financial. Macro-level evaluation computes precision, recall, and F1 scores independently for each BIO-concept label derived from the US-GAAP taxonomy, and then averages them without weighting. This ensures that each concept, including rare or infrequent ones, contributes equally to the final score, making it especially suitable for domains with skewed label distributions. In contrast, micro-level evaluation aggregates token-level true positives, false positives, and false negatives across all labels before computing precision, recall, and F1. This approach emphasizes the model’s overall tagging accuracy by treating every token equally and thus better reflects performance on frequent concepts. Together, these two metrics provide a balanced view of both per-concept performance and overall tagging quality.
@@ Appendix E Evaluation Metrics
$\text{Precision}=\frac{|\mathcal{G}\cap\mathcal{P}|}{|\mathcal{P}|}$ (8)
$\text{Recall}=\frac{|\mathcal{G}\cap\mathcal{P}|}{|\mathcal{G}|}$ (9)
$\text{F1}=\frac{2\times\text{Precision}\times\text{Recall}}{\text{Precision}+\text{Recall}}$ (10)
@@
$\text{Accuracy}=\frac{1}{N}\sum_{i=1}^{N}\delta(\hat{c}_{i},c_{i}^{*})$ (11)
@@
Appendix F Evaluation Models Details